OpenStudy (anonymous):

I really need help with understanding this problem Find a formula for the quadratic function with the points (2,0), (-1, 0), and (0,4)

OpenStudy (anonymous):

so far what I have is 4a+2b+c = 0 a-b+c=0 c=4 -4a-2b-c=0 a-b+c=0 3(-3a-3b=4) -a+b-c=0 c=4 -a+b=4 -9a-9b=12 -a+b=4 -10a-8b=16 but the b does not cancel out to give me a

OpenStudy (anonymous):

@mathslover @dan815 @Hero @jim_thompson5910 @SolomonZelman

OpenStudy (anonymous):

I set up three equations from the points and then I added equation 1 to equation 2 and equation 2 to equation 3.

OpenStudy (dan815):

y=ax^2+bx+c c=4 from point (0,4) a-b+4=0 b=a+4 from point(-1,0) 4a+2b+4=0 sub b=a+4 4a+2(a+4)+4=0 4a+2a+8+4=0 a=-12/6 = -2 and b=a+4 = -2+4 = 2 therfore y=-2x^2+2b+4

OpenStudy (anonymous):

what did you do in the second step?

OpenStudy (dan815):

which one

OpenStudy (anonymous):

so I understand where you got the value for c but a-b+4=0 b = a+4

OpenStudy (anonymous):

are you just substituting the value of c in one of the equations?

OpenStudy (anonymous):

and the 4a+2b+4 = 0 where is this from?

OpenStudy (anonymous):

@dan815